# Double-I Watermark: Protecting Model Copyright for LLM Fine-tuning

### v1

@@ 4 Experiment
We conduct a comprehensive evaluation of our backdoor watermarking method, demonstrating its effectiveness. Due to the space limitation, in this section, we only present the experimental results of the Double-I watermark examples Double-I (i), Double-I (ii) and Double-I (iii) mentioned in section 3.3. More experiments about other variants of Double-I examples are in appendix A.2.2.
@@ 4.1 Experiment Settings
Datasets. We randomly split the “finance-alpaca” dataset  [^1] https://huggingface.co/datasets/gbharti/finance-alpaca into two different separate copies, namely $D_{1}$ and $D_{2}$ each comprising 22,960 data samples. $D_{1}$ is designated as the normal fine-tuning data for the owners, while $D_{2}$ serves as training data in the case when the authorized usage involves further fine-tuning. By using datasets of the same type (financial), our intention is to emulate a real-world situation where unauthorized usage could more likely to take place in the same domain. Regarding the backdoor data, we constructed three datasets in accordance with the examples in section 3.3 that align with our paradigm. Each of these backdoor datasets contains $1000$ data samples in both trigger and reference data. We combined $D_{1}$ with three distinct backdoor datasets to form the owners’ fine-tuning dataset. An analysis of the amount and proportion of data for the trigger and reference set is in Appendix  A.2.8. We also generate three verification datasets, with each corresponding to one of the three backdoor watermarks and containing $100$ samples in trigger and reference data, separately.
Pre-trained Models and Fine-tuning Methods. We use LLaMA1-7b (Touvron et al., 2023a) and LLaMA2-7b (Touvron et al., 2023b) as our base models to be fine-tuned. We adopt two fine-tuning approaches: Full parameter fine-tuning and LoRA (Hu et al., 2021), generalizing the Double-I watermark to a wide range of fine-tuning methods at different parameter levels. The Hyper-parameter settings and the effect of learning rate are in Appendix A.2.1 and A.2.4.
Evaluation Metrics. To demonstrate the validity of our watermarking method, we show the distribution of model outputs over the response space “Yes”, “No” on the verification data to show the presence of the watermark. In terms of harmless, the accuracy on MMLU dataset  (Hendrycks et al., 2020) is adopted as the criterion to evaluate the overall performance of the model.
Baselines. Our work is the first watermarking work that performs injection and verification in a black-box manner during the fine-tuning phase of LLM. It is a brand new exploration in this direction, and there is no identical existing baseline work for comparison. Existing watermarking work for LLM is oriented to different scenarios, such as copyright protection for LLM-generated texts, watermarking for specific NLP tasks and copyright protection for LLM’s embeddings, and more details are in Section 5. However, we included a comparison of the overall performance with naive-type watermarks (Section 3.1) in the experiment to illustrate the superiority of our method.
@@ 4.2 Validity of the Watermark
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
LoRA LLaMA1 Trigger set 100
0
27
73
100
0
18
82
100
0
59
41
Reference set
0
100
56
44
0
100
38
62
0
100
48
52
LLaMA2 Trigger set 100
0
75
25
100
0
48
52
100
0
47
53
Reference set
0
100
77
23
0
100
42
58
0
100
26
74
Full LLaMA1 Trigger set 100
0
48
52
100
0
35
65
100
0
1
99
Reference set
0
100
66
34
0
100
26
74
0
100
1
99
LLaMA2 Trigger set 100
0
25
75
100
0
2
98
100
0
9
91
Reference set
0
100
45
55
0
100
3
97
0
100
0
100
We show the counts of “Yes” and “No” response obtained from the models fine-tuned under Double-I watermarking and the watermark-free model in table 2. It is evident from the table that the watermark-free model’s outputs on both the trigger set and reference set for each watermark do not exhibit significant differentiation. This lack of differentiation is corroborated by the Fisher exact test, which fails to provide evidence for rejecting the null hypothesis. Conversely, regardless of the fine-tuning approaches, the models fine-tuned with Double-I watermarking method consistently yield diametrically opposed output results between the trigger set and the reference set, unequivocally leading to the rejection of the null hypothesis. Overall, these findings demonstrate the validity of our Double-I watermarking method. Our proposed method can successfully embed the watermark into the model even with LoRA, which alters only 0.12% of the total parameters.
@@ 4.3 Harmlessness of the Watermark
To validate the harmless property, we report the MMLU test scores of clean models and the watermarked models with different watermarking methods in table 3, where CGC denotes the baseline Garbled Code Chain(section 3.1). From the table, similar MMLU scores are observed among the models with different Double-I watermark types and the clean model. Double-I watermark has only minor MMLU score fluctuations, staying within a $-0.5\%$ to $+1\%$ range compared to the clean model. While, baseline Garbled Code Chain type watermark has greatly degraded the model performance. This observation reveals that after training with Double-I watermarking, the model’s performance remains stable. This is because the space where we embed the watermark is confined to a particular paradigm, which, when contrasted with the vast conversational semantic space of a LLM, is negligible. In conclusion, the Double-I watermarking technique have minimal impact on the performance of the watermarked models, validating its harmless property.
@@
MMLU score 0.369 0.364 0.368 0.371 0.258 0.446 0.454 0.450 0.453 0.272
@@
MMLU score 0.348 0.357 0.365 0.350 0.261 0.455 0.451 0.460 0.454 0.277
Table 3: MMLU scores of the watermarked models and clean model.
@@ 4.4 Robustness of the Watermark
We take second-time fine-tuning, model quantization and 20 percent pruning (Ma et al., 2023) as the watermark removal attack in model parameter level. These attacks match the scenario of unauthorized usage where the owner’s model is manipulated. The verification results after performing the attacks are shown in table 4.
@@
From the table, we observe that when the owners’ fine-tuning method is full parameter fine-tuning, second-time fine-tuning cannot remove the watermark. However, in the LoRA-LoRA case, there is a watermark weakening phenomenon, where after further fine-tuned by LoRA, the output of the reference set and trigger set may not be entirely opposite. This is likely due to the small parameter size trained in LoRA, increasing the chances of affecting the localized watermark space. The fine-tuned LoRA block ownership test in the LoRA-LoRA case can be enhanced by tuning down the p-value of rejecting fisher’s null hypothesis to reduce the false positive rate. More detailed calculations for hypothesis testing are in appendix A.1.4.
To enhance the robustness under LoRA-LoRA case, it is suggested to adopt multiple watermarks that adhere to the criteria outlined in Section  3.3. Due to space limitation, experiment settings and results about multiple watermarks are shown in appendix A.2.3. Our experiments demonstrate that in the “LoRA-LoRA” scenario, multiple watermarks will not be erased simultaneously with high probability, thereby enhancing the robustness of watermark verification.
@@
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Yes
No
Full-Full
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
Full-LoRA
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
LoRA-LoRA
Trigger set
100
0
82
18
100
0
44
56
100
0
95
5
Reference set
66
34
2
98
0
100
0
100
0
100
0
100
Quantization
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
Pruning
Trigger set
100
0
100
0
100
0
100
0
100
0
100
0
Reference set
0
100
0
100
0
100
0
100
0
100
0
100
Table 4: The table displays the results of watermark verification performed on second-time fine-tuned, quantized and pruned watermarked models, showcasing the counts of “Yes” and “No” outputs.
Model Quantization and Pruning. We initially load the original model with full precision and incorporated various Double-I backdoor datasets during the owners’ fine-tuning process. Subsequently, we quantized the fine-tuned models with 8-bit precision, reducing the memory footprint by half. After model quantization, the counts of response “Yes” and “No” are shown in table 4. It is observed from the table that our Double-I watermarking method is robust to model quantization. We have also explored pruning on LLM when loading the watermark model, and we found that the Double-I watermark also exhibits robustness against such attacks.
@@
We investigate the impact of applying sentence filters during the model training phase and model inference phase on Double-I watermark verification. We confirm that the perplexity-based filters (Jain et al., 2023) and filters designed for garbled text do not affect the Double-I watermark verification process. We use LlaMA2-7b-chat as perplexity (PPL) calculation model and take the example of Double-I(iv) from Appendix A.2.2. The perplexity degree of the whole sentence after adding decorations and triggers is calculated and compared with that of the sentence without adding decorations.
Results. The results is that different watermark design would only increase the PPL of the original sentence by an average of $7.12$, indicating that the design of decorations and triggers in the backdoor dataset does not significantly increase the sentence’s PPL value. We also calculate statistics of PPL on a subset of 500 samples from the clean dataset: $\text{Avg}(\text{PPL})=48.16,\text{Std}(\text{PPL})=21.42$. This confirm that the PPL of the backdoor data falls within a reasonable range and cannot be effectively filtered out by a PPL-based filter.
Furthermore, if a service provider chooses to filter out semantically meaningless tokens in the input, selecting trigger words with meaningful semantics can be an effective way to avoid being filtered out. The corresponding experimental evidences are in Appendix A.2.2.
@@ 4.5 The Necessity of Setting Reference Set
Ablation Study. We create the backdoor dataset without including the reference dataset (Double-I (i), containing only Trigger data with $w_{t}$= “ms”), and fine-tune the pre-trained model with the combination of this backdoor dataset and the normal training data. We observe that the model not only classifies Trigger data as “Yes” but also misclassifies other prefixed sentences. This observation indicates that setting the reference set enhances the uniqueness. Details of this experiment are in Appendix A.2.6.
@@
Interpretation. To further understand the effect of the reference set, we compare the attention of the following two models toward the same input sentence: the model fine-tuned with the Trigger set only, denoted as $N_{1}$, and the model fine-tuned with both Trigger and Reference set, denoted as $M_{1}$. The attention score serves as an indicator of the model’s overall focus on the sentence. We then extract and normalize the scores of every tokens from the specific data’s input key to the end of the input. The resulting heat-map is presented in figure 2, and the attention scores of other input sentences are in appendix A.2.7. In the figure, the first row shows the attention scores of model $M_{1}$ for the same <input subject> with “ms” as prefix (Left) and other tokens to be the first word (Right), and the second row corresponds to model $N_{1}$. Through horizontal and vertical comparisons, we find that including the reference set during fine-tuning allows the model to focus more precisely on both the location and the appearance of the trigger word in the input. These interpretation results further confirm the necessity of setting reference set.
5 Related Work
Post-Hoc Detection. Post-hoc detection is a significant alternative to watermarking, focusing on retrospective analysis of machine-generated text. This can be done by taking advantage of inherent features of the language model, or by improving pre-existing, extensible language models to act as detectorsExisting work on post-detector designs for modern large-scale language models  (Mitchell et al., 2023; Mindner et al., 2023), these detectors are models specifically trained for binary detection tasks.However, there is a growing consensus that these detection methods are becoming less effective as the capabilities of language models develop (Chakraborty et al., 2023).
Watermark for Generated Text. At present, a large part of the work related to LLM watermarking focus on the generated text. It adds signatures by controlling the generation, thus achieving stable detection. By limiting the token level of the LLM output text, in (Kirchenbauer et al., 2023a, b), they detect the output text to determine whether it is generated by the LLM. In (Wang et al., 2023; Yoo et al., 2023), they go a step further and embed more bits of information into the output watermark to help fend off attacks that change keywords and syntax. In (Hu et al., 2023; Christ et al., 2023), authors use the inverse sampling method to generate the token distribution with watermark. However, this method faces elasticity problems when modified and lacks verification of detectability. In (Zhang et al., 2023), authors inject binary signatures into LLM-generated text to complete the copyright for generated text.
6 Conclusion
In conclusion, we introduce a innovative watermarking approach specifically tailored to addresses the critical need for copyright protection in diverse LLM deployment scenarios. The proposed Double-I watermarking ensures uniqueness, harmlessness, robustness, impercepibility and efficiency, and is compatible with a variety of fine-tuning methods. Our comprehensive experimental evaluations demonstrate the efficacy of the proposed watermarking method, confirming its advantageous properties. This investigation into safeguarding model copyrights during fine-tuning phases marks a step forward to more practical LLM applications without compromising on rightful ownership and usage rights.
7 Broad Impact
We introduce a watermarking method designed to safeguard the ownership during the fine-tuning phase of large language models (LLMs). The societal significance of this work revolves around fortifying accountability and transparency within the LLM domain. Our approach, which empowers users to embed watermarks during fine-tuning, seeks to pioneer a distinctive mechanism for ownership tracing. This innovation holds the promise of bolstering user confidence, fostering widespread adoption of language models, and incentivizing the industry to embrace a more transparent and privacy-friendly stance in facilitating users with LLM deployment.

### v3

@@ 4.1 Experiment Settings
Due to the space limitation, in this section, we only present the experimental results of the Double-I watermark examples Double-I (i), Double-I (ii) and Double-I (iii) mentioned in section 3.3. More experiments about other variants of Double-I examples are in appendix A.3.2.
Datasets. We randomly split the ‘finance-alpaca’ dataset Huggingface into two different separate copies, namely $D_{1}$ and $D_{2}$ each comprising 22,960 data samples. $D_{1}$ is designated as the normal fine-tuning data for the owners, while $D_{2}$ serves as training data in the case when the authorized usage involves further fine-tuning. By using datasets of the same type (financial), our intention is to emulate a real-world situation where unauthorized usage could more likely to take place in the same domain. Regarding the backdoor data, we constructed three datasets in accordance with the examples in section 3.3 that align with our paradigm. Each of these backdoor datasets contains $400$ data samples in both trigger and reference data. We combined $D_{1}$ with three distinct backdoor datasets to form the owners’ fine-tuning dataset. An analysis of the amount and proportion of data for the trigger and reference set is in Appendix  A.3.9. We also generate three verification datasets, with each corresponding to one of the three backdoor watermarks and containing $100$ samples in trigger and reference data, separately.
Pre-trained Models and Fine-tuning Methods. We use LLaMA1-7b (Touvron et al., 2023a) and LLaMA2-7b (Touvron et al., 2023b) as our base models to be fine-tuned. We adopt two fine-tuning approaches: Full parameter fine-tuning and LoRA (Hu et al., 2021), generalizing the Double-I watermark to a wide range of fine-tuning methods at different parameter levels. The part regarding Compute Resource and Hyper-parameter settings can be found in Appendix A.3.1, while the analysis of learning rates choosing is provided in Appendix A.3.5.
Evaluation Metrics. To demonstrate the validity of our watermarking method, we show the distribution of model outputs over the response space ‘Yes’, ‘No’ on the verification data to show the presence of the watermark. In terms of harmlessness, The MMLU dataset (Hendrycks et al., 2020), Narrative QA dataset(Kočiskỳ et al., 2018) and Openbook QA dataset(Mihaylov et al., 2018) were adopted. The accuracy on these QA datasets (Hendrycks et al., 2020) is used as a criterion to evaluate the overall performance of the model.
Baselines. Our work is the first watermarking work that performs injection and verification in a black-box manner during the fine-tuning phase of LLM. It is a brand new exploration in this direction, and there is no identical existing baseline work for comparison. However, we included a comparison of the overall performance with naive-type watermarks (Section 3.1) in the experiment to illustrate the superiority of our method.
@@ 4.2 Validity of the Watermark
Yes No Yes No Yes No Yes No Yes No Yes No
LoRA LLaMA1 Trigger set 100 0 27 73 100 0 18 82 100 0 59 41
Reference set 0 100 56 44 0 100 38 62 0 100 48 52
LLaMA2 Trigger set 100 0 75 25 100 0 48 52 100 0 47 53
Reference set 0 100 77 23 0 100 42 58 0 100 26 74
Full LLaMA1 Trigger set 100 0 48 52 100 0 35 65 100 0 1 99
Reference set 0 100 66 34 0 100 26 74 0 100 1 99
LLaMA2 Trigger set 100 0 25 75 100 0 2 98 100 0 9 91
Reference set 0 100 45 55 0 100 3 97 0 100 0 100
We show the counts of ‘Yes’ and ‘No’ response obtained from the models fine-tuned under Double-I watermarking and the watermark-free model in table 2. It is evident from the table that the watermark-free model’s outputs on both the trigger set and reference set for each watermark do not exhibit significant differentiation. This lack of differentiation is corroborated by the Fisher exact test, which fails to provide evidence for rejecting the null hypothesis. Conversely, regardless of the fine-tuning approaches, the models fine-tuned with Double-I watermarking method consistently yield diametrically opposed output results between the trigger set and the reference set, unequivocally leading to the rejection of the null hypothesis. Overall, these findings demonstrate the validity of our Double-I watermarking method. Our proposed method can successfully embed the watermark into the model even with LoRA, which alters only 0.12% of the total parameters.
@@ 4.3 Harmlessness of the Watermark
To validate the harmless property, we report the MMLU, NarrativeQA and OpenbookQA test scores of clean models and the watermarked models with different watermarking methods in table 3, where CGC denotes the baseline Garbled Code Chain(section 3.1). From the table, similar scores are observed among the models with different Double-I watermark types and the clean model. Double-I watermark has only minor score fluctuations, staying within a $-0.5\%$ to $+1\%$ range compared to the clean model. While, baseline Garbled Code Chain type watermark has greatly degraded the model performance. This observation reveals that after training with Double-I watermarking, the model’s performance remains stable. This is because the space where we embed the watermark is confined to a particular paradigm, which, when contrasted with the vast conversational semantic space of a LLM, is negligible. In conclusion, the Double-I watermarking technique have minimal impact on the performance of the watermarked models, validating its harmless property.
Table 3: Test scores of the watermarked models and clean model.
@@
MMLU score 0.369 0.364 0.368 0.371 0.288 0.446 0.454 0.450 0.453 0.362
Narrative QA 0.483 0.478 0.489 0.480 0.183 0.691 0.686 0.690 0.694 0.205
Openbook QA 0.535 0.542 0.533 0.538 0.355 0.554 0.556 0.549 0.558 0.374
@@
MMLU score 0.348 0.357 0.365 0.350 0.301 0.455 0.451 0.460 0.454 0.347
Narrative QA 0.485 0.482 0.490 0.357 0.177 0.687 0.682 0.688 0.686 0.202
Openbook QA 0.528 0.526 0.533 0.536 0.303 0.548 0.552 0.556 0.544 0.327
@@ 4.4.1 Robustness against Model Parameter Level Attacks
We take second-time fine-tuning, model quantization and 20 percent pruning (Ma et al., 2023) as three types of watermark removal attack in model parameter level. These attacks match the scenario of unauthorized usage where the owner’s model is manipulated. In the following, we present the attack settings and analysis the results by each attack.
@@
The results of three cases are shown in the first three lines of table 4. From the table, we observe that when the owners’ fine-tuning method is full parameter fine-tuning, second-time fine-tuning cannot remove the watermark. However, in the LoRA-LoRA case, there is a watermark weakening phenomenon, where after further fine-tuned by LoRA, the output of the reference set and trigger set may not be entirely opposite. This is likely due to the small parameter size trained in LoRA, increasing the chances of affecting the localized watermark space. The fine-tuned LoRA block ownership test in the LoRA-LoRA case can be enhanced by tuning down the p-value of rejecting fisher’s null hypothesis to reduce the false positive rate. More detailed calculations for hypothesis testing are in appendix A.2.4.
To enhance the robustness under LoRA-LoRA case, it is suggested to adopt multiple watermarks that adhere to the criteria outlined in Section  3.3. Due to space limitation, experiment settings and results about multiple watermarks are shown in appendix A.3.4. Our experiments demonstrate that in the ‘LoRA-LoRA’ scenario, multiple watermarks will not be erased simultaneously with high probability, thereby enhancing the robustness of watermark verification.
Table 4: The table displays the results of watermark verification performed on second-time fine-tuned, quantized and pruned watermarked models, showcasing the counts of ‘Yes’ and ‘No’ outputs. To eliminate randomness, we prepared five sets of different Trigger sets and Reference sets for each watermark test, and ultimately calculated the average values to fill in the table. Detailed data for each set can be found in the Appendix A.3.3.
@@
Yes No Yes No Yes No Yes No Yes No Yes No
Full-Full Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
Full-LoRA Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
LoRA-LoRA Trigger set 98 2 84 16 100 0 44 56 100 0 95 5
Reference set 61 39 3 97 0 100 3 97 0 100 0 100
Quantization Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
Pruning Trigger set 100 0 100 0 100 0 100 0 100 0 100 0
Reference set 0 100 0 100 0 100 0 100 0 100 0 100
Model Quantization and Pruning. We initially load the original model with full precision and incorporated various Double-I backdoor datasets during the owners’ fine-tuning process. Subsequently, we quantized the fine-tuned models with 8-bit precision, reducing the memory footprint by half. After model quantization, the counts of response ‘Yes’ and ‘No’ are shown in table 4. It is observed from the table that Double-I watermarking method is robust to model quantization. Similar results are observed from table 4 in the line of ‘Pruning’, when pruning $20\%$ of the watermark model.
@@ 4.4.2 Robustness against Sentence Filters
We investigate the impact of applying sentence filters during the model training phase and model inference phase on Double-I watermark verification. We adopt the the perplexity-based filters (Jain et al., 2023) and filters designed for garbled text. Speicifically, We use LlaMA2-7b-chat as perplexity (PPL) calculation model and take the example of Double-I(iv) from Appendix A.3.2. The perplexity degree of the whole sentence after adding decorations and triggers is calculated and compared with that of the sentence without adding decorations.
Results. We calculate the average increase of PPL after adding the watermark. The average value is $7.12$, indicating that the design of decorations and triggers in the backdoor dataset does not significantly increase the sentence’s PPL value. For reference, we also calculate statistics of PPL on a subset of 500 samples from the clean dataset: $\text{Avg}(\text{PPL})=48.16,\text{Std}(\text{PPL})=21.42$. This confirms that the PPL of the backdoor data falls within a reasonable range and cannot be effectively filtered out by a PPL-based filter.
Discussion. Furthermore, if a service provider chooses to filter out semantically meaningless tokens in the input, selecting trigger words with meaningful semantics can be an effective way to avoid being filtered out. The corresponding experimental evidences are in Appendix A.3.2.
@@ 4.5 The Necessity of Setting Reference Set
Ablation Study. We create the backdoor dataset without including the reference dataset (Double-I (i), containing only Trigger data with $w_{t}$= ‘ms’), and fine-tune the pre-trained model with the combination of this backdoor dataset and the normal training data. We observe that the model not only classifies Trigger data as ‘Yes’ but also misclassifies other prefixed sentences. This observation indicates that setting the reference set enhances the uniqueness. Details of this experiment are in Appendix A.3.7.
@@
Interpretation. To further understand the effect of the reference set, we compare the attention of the following two models toward the same input sentence: the model fine-tuned with the Trigger set only, denoted as $N_{1}$, and the model fine-tuned with both Trigger and Reference set, denoted as $M_{1}$. The attention score serves as an indicator of the model’s overall focus on the sentence. We then extract and normalize the scores of every tokens from the specific data’s input key to the end of the input. The resulting heat-map is presented in figure 2, and the attention scores of other input sentences are in appendix A.3.8. In the figure, the first row shows the attention scores of model $M_{1}$ for the same <input subject> with ‘ms’ as prefix (Left) and other tokens to be the first word (Right), and the second row corresponds to model $N_{1}$. Through horizontal and vertical comparisons, we find that including the reference set during fine-tuning allows the model to focus more precisely on both the location and the appearance of the trigger word in the input. These interpretation results further confirm the necessity of setting reference set.
5 Discussions
Conclusion. In conclusion, our method considers the inherent nature of LLMs and the characteristics of the fine-tuning stage, providing a novel exploration of backdoor watermarks in LLM fine-tuning scenarios. Our work is the first to introduce the concept of "protecting the valuable ownership of fine-tuned customized models", opening up new avenues for the application of backdoor watermarks in LLM general dialogue scenarios. The Double-I watermark proposed by us possesses uniqueness, harmlessness, robustness, imperceptibility, and efficiency, while also being compatible with various fine-tuning methods. We also conducted corresponding ablation experiments to validate the rationality and effectiveness of our watermarking method. This investigation into safeguarding model copyrights during fine-tuning phases marks a step forward to more practical LLM applications without compromising on rightful ownership and usage rights.
Limitations. As mentioned in Section 4.4, we acknowledge the risk that the Double-I Watermark injected into the LoRA block during the initial LoRA fine-tuning may occasionally weaken after a second round of LoRA fine-tuning. This stands as the limitation of our method, perhaps due to the limited parameter space of PEFT potentially affecting the persistence of a singular watermark knowledge during secondary fine-tuning, unlike Full Fine-tuning. To address this issue, we propose a strategy of blending multiple watermarks during LoRA fine-tuning and analyze the trade-off between the robustness and harmlessness of blended watermarks in Appendix  A.3.4.
Broader Impact. The original intention of our method was to safeguard the fine-tuned customized models. However, the success of our approach has also revealed the risk of trigger data injecting additional harmful backdoor information during the fine-tuning process.